# A Kalman Filter for track reconstruction in very large time projection chambers

### v1

@@ Abstract
This study introduces a Kalman Filter tailored for homogeneous gas Time Projection Chambers (TPCs), adapted from the algorithm utilized by the ALICE experiment. In order to describe semi-circular paths in the plane perpendicular to the magnetic field, we introduce a novel mirror rotation technique into the Kalman Filter algorithm, enabling effective tracking of trajectories of varying lengths, including those with multiple circular paths within the detector, also known as “loopers”. Demonstrated relative improvements of up to 80% in electron momentum resolution and up to 50% in muon and pion momentum resolution underscore the significance of this enhancement. Such advancements hold promise not only for the future of the ALICE TPC but also for neutrino high-pressure gas TPCs, where loopers become significant owing to the randomness of production points and their relatively low energies in neutrino interactions.
@@ 2 The Kalman Filter
    Update: Produce an updated estimate for the state vector and the covariance matrix, $s_{k}$ and $C_{k}$, respectively, using information from the current measurement. The update is performed, requiring that the covariance for the new estimate is minimized. The update transformation is based on the use of the so called Kalman gain:
@@
    Using the Kalman gain, the update operation, also known as filtering, follows as:
@@ 2.1 The custom Kalman Filter
Multiple scattering is treated through the noise correction matrix, $Q_{k}$. At each step the scattering angle, $\theta_{\textrm{M}}$, is calculated using the formula given by Lynch and Dahl [37]:
@@
Each step in the evolution of the Kalman Filter can potentially fail, in which case the algorithm is stopped. This can happen mainly in two scenarios: $\sin\phi$ can be calculated to be out of range, i.e. $|\sin\phi|>(1-10^{-7})$ or the particle can lose all its remaining energy. Once the Kalman Filter is stopped, the information for each of the reconstructed points is saved, while only the truth level information is recorded for the other points. Flags are used to preserve information on which of the reconstruction steps have been successful and which have failed.
@@
Figure 3: Flow chart describing the CKF algorithm. A seeding algorithm is used to obtain an estimate for the status of the system at the start of the trajectory $k=0$. Energy loss and multiple scattering corrections are applied to the estimate. The fit is then moved to the next point $k\rightarrow k+1$ either by applying the the BKF procedure or by using the mirror rotation, in the case that the limits of the $\sin\phi$ range have been surpassed. The algorithm is iterated point by point until the end of the trajectory is reached.
@@ 3 Toy Monte Carlo Simulation
To generate particle samples and validate the CKF algorithm, we employed a toy Monte Carlo (MC) tool called fastMCKalman [31]. This tool, stemming from the AliExternalTrackParam framework in the AliRoot code-base [28], was designed to be complemented by RootInteractive [32], an advanced statistical analysis tool. fastMCKalman has been developed with several objectives in mind: conducting rapid Monte Carlo (MC) simulations to evaluate tracking performance metrics, particle identification, and time-of-flight measurements across various detector setups. It was also designed to facilitate detailed studies on signal distortion in the ALICE detector and the derivation of performance metrics for its Run-3 upgrade and future iterations. In this discussion, we highlight the effectiveness of fastMCKalman in rapid simulation and tracking performance assessment within a TPC setting.
@@
The propagation of the state vector is obtained by applying Eq. 15. At each step the energy loss is calculated using the Bethe-Bloch formula as described in Eq. 16 and multiplying by the distance traveled and the material density. The energy loss is then converted in the $q/{p_{\text{T}}}$ multiplicative factor described in Eq. 15 and smeared with a Landau distribution having a width equal to the $c_{k}$ factor described in Eq. 17. The multiple scattering effects are simulated by calculating the scattering angle $\theta_{\textrm{M}}$ and the process noise matrix $Q$. The diagonal elements of the matrix are then used as the widths of smearing Gaussian distributions that are applied to parameters $s_{2}=\sin\phi,s_{3}=\tan\lambda$ and $s_{4}=q/p_{\text{T}}$. To reproduce the measurement noise encapsulated in matrix $R$, a Gaussian smearing is applied to the position parameters $s_{0}=y$ and $s_{1}=z$. The widths of the distributions are equal to the position resolutions $\sigma_{r\phi}$ and $\sigma_{z}$ respectively.
The propagation continues until any of the following happens: the particle reaches the edges of the detector cylinder; the energy loss of the particle gets so close to the Bragg peak region that the Bethe-Bloch model is no longer applicable; the track has traversed a predefined maximum number of points; one of the propagation steps fails. Once the track is fully generated, the track fit is done as described in Sec. 2.1. No element of track formation or particle identification is included. After saving all information, a new simulation and track fit start.
@@ 3.1 Sample definition
The tracking pad response as well as the gas properties of the detector were sampled in each simulated event: The resolutions $\sigma_{r\phi}=\sigma_{z}$ were uniformly distributed between 0.1 cm and 0.5 cm and the pressure $P_{\textrm{gas}}$ was randomized between 0.1 atm and 10 atm. The gas composition was taken to be the Ne/CO2/N2 (90/10/5) gas mixture used by the ALICE experiment during Run-1 [6]. The radiation length and density of the gas at atmospheric pressure are $X_{0}=1.2763\times 10^{4}\ \text{cm}$ and $\rho=0.0016265\ \text{g/cm}^{3}$. The particles produced are equally divided in electrons, muons, pions, kaons and protons, corresponding to the ALICE’s convention for particle types, $t_{\textrm{ID}}$, 0, 1, 2, 3, and 4, respectively. The angles $\phi$ and $\lambda$ are fully randomized. The initial $p_{\textrm{T}}$ is sampled from a two-component distribution: a high-$p_{\text{T}}$ component uniformly distributed in $[0,20]$ GeV/c, which covers 70% of the total, and a low-$p_{\text{T}}$ component flat in $1/p_{\textrm{T}}$:
@@
where $p_{\textrm{T}_{\textrm{min}}}=0.01$ GeV$/c$, $p_{\textrm{T}_{\textrm{max}}}=20\ \text{GeV}/c$ and $j$ is a random variable uniformly distributed between 0 and 1. Some key properties of the tracks composing the PS sample are plotted in Fig. 5. These include the $p_{\textrm{T}}$ spectrum, the lever arm $L_{\textrm{Arm}}$ and the number of points per track $N$ separated between primaries and secondaries. The lever arm is defined as the distance in the $xy$ plane between the first and last point in the track. All the tracks included in this and future plots have been successfully reconstructed, unless stated otherwise. The PS sample track reconstruction efficiencies $\epsilon$ as a function of the same variables considered in Fig. 5 are shown in Fig. 19 in the Appendix.
@@
Figure 5: Distributions of (a) transverse momentum $p_{\textrm{T}}$, (b) lever arm $L_{\textrm{Arm}}$ and (c) number of points per track $N$ in the PS sample. In all the plots the distributions for primary and secondary particles are shown separately. A at low $p_{\textrm{T}}$ is produced by the sample component regulated by Eq. 28. The spikes around $N=250$ and $L_{\textrm{Arm}}=250$ in the primary sample are explained by the simulated geometry having 250 radial pad layers and the particles starting from the center of the detector.
Primaries and secondaries have analogous $p_{\textrm{T}}$, but significant differences in their $L_{\textrm{Arm}}$ and $N$. Since the primaries all start at the center of the detector, most tracks will cross the detector exiting from the cylinder’s barrel, producing a track with as many points as the 250 pad layers. In alternative the track can exit from the sides of the detector, producing tracks with a slightly smaller $N$ or be stopped inside the detector having $N<50$. For the secondaries the spread is much more homogeneous and the chance of producing tracks with $N>250$ is more significant.
@@
A second sample containing a total of $10^{5}$ particle tracks was produced to recreate conditions analogous to the ones that would be experienced by a HPgTPC in a accelerator neutrino experiment, such as the ND-GAr detector. The goal for this second sample is to explore the potential performance of such a detector, using realistic particle spectra and spatial resolutions. This sample will be referred to as the high-pressure sample or HP sample. The HP sample is produced with randomized starting positions in the same manner as the ones applied to the secondaries in the PS sample. This is done to emulate the randomness of particle track formation in a neutrino experiment. The detector characteristics are fixed, having the same cylinder dimensions and pad distribution of the previous sample. The point resolutions are taken as $\sigma_{r\phi}=\sigma_{z}=0.1\ \text{cm}$, comparable to what is quoted in ALICE [15]. The gas is a mixture of argon and methane at a 90 to 10 ratio at 10 atm of pressure, which is the nominal gas suggested for the ND-GAr detector in the DUNE Near Detector [13]. This composition corresponds to a $X_{0}=1.193\times 10^{3}\ \text{cm}$ and a density of $\rho=0.01677\ \text{g/cm}^{3}$. Only three particle types were considered in this case: muons, pions and protons. These were chosen because they are the key particles produced in $\nu_{\mu}$ charged-current interactions that are the most relevant in an accelerator neutrino experiment such as DUNE. The initial transverse momenta were randomized to be uniformly distributed between 0.01 GeV$/c$ and 5 GeV$/c$ and the angles are randomized over the whole spectrum. The $p_{\text{T}}$, $L_{\textrm{Arm}}$ and $N$ distributions for the sample are shown in Fig. 6. Similar to the PS sample, the reconstruction efficiencies are shown in the Appendix in Fig. 20.
@@ 3.2 Tests and results: parameter scan sample
The pulls were tested for the sample, both for the results of the Seed and for the estimates evaluated at the start of the track, after the full propagation of the CKF: the resulting distributions for all the state vector parameters are shown in Figs. 7 and 8, respectively. All the pull distributions were fitted to a standard Gaussian distribution and were found to be centered in 0 and have $\sigma\sim 1$; this implies that the diagonal elements of the covariance matrices well describe the uncertainties.
@@
Figure 8: Pull distributions obtained after the full propagation of the CKF algorithm over the whole PS sample. See Fig. 7 for comparison.
@@
where in our case, $\mu$ corresponds to the true value of state vector, $s^{\text{true}}$, $s$ and $C$ are the estimates obtained from the reconstruction and $n=5$. The Mahalanobis distance, $M$, of a set of points belonging to the distribution $D$, follows a $\chi^{2}$ distribution with $n$ degrees of freedom. One can check if $C$ is well defined, by verifying that the corresponding $M$ follow a $\chi^{2}$ distribution with the correct number of degrees of freedom. In Fig. 9, we show the results of a $\chi^{2}$ fit over the $M$ distribution for the whole sample. The plot on the left shows the results obtained from the Seed algorithm, while the one on the right shows the results after the full CKF propagation. In both cases the n.d.f. obtained with the $\chi^{2}$ fit are very close to $n=5$: both Seed and CKF estimates accurately the covariances of the reconstructed track parameters in the state vector.
Figure 9: Mahalanobis distance $M$ distribution for the PS Sample fitted by a standard $\chi^{2}$ p.d.f. showing the results for the $n.d.f.$ parameter. The expected result for a 5-dimentional matrix is $n.d.f.=5$. The results for the Seed and for the fully propagated CKF are shown in plots (a) and (b) respectively.
@@
where $l$ is the length of the track in the $xy$-plane. Note that the value of $1/\left(\beta p_{\text{T}}\right)$ is averaged along the trajectory to take into account energy loss. In Figs. 1011 and 12, the upper plots show the CKF covariance estimates, $\sqrt{C_{44}^{\textrm{CKF}}}$, while the bottom plots show their ratios to the theoretical expectations, $\sqrt{C_{44}^{\textrm{theo}}}$. The points analyzed are randomly taken along the reconstructed tracks and down-sampled to $10\%$ of the total to avoid correlations.
@@
Figure 10: (a) CKF $q/p_{\text{T}}$ resolution $\sigma_{\text{kf}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{CKF}}}$ as a function of the true $q/p_{\text{T}}$. (b) Ratio of the CKF $q/p_{\text{T}}$ resolution, over the theoretical expectations $\sigma_{\text{theo}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{theo}}}$, as a function of the true $q/p_{\text{T}}$. The histograms include all particles in the PS sample and are color-coded according to the ALICE convention for particle types $t_{\textrm{ID}}=(0,1,2,3,4)=(\textrm{e},\mu,\pi,K,\textrm{p})$. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical.
Figure 11: Similar plots to the ones presented in Fig. 10. The histograms in this case are color-coded according to the according to the gas pressure $P_{\textrm{gas}}$.
Figure 12: Similar plots to the ones presented in Fig. 10. The histograms in this case are color-coded according to the according to the radial resolution $\sigma_{r\phi}$.
@@
In Fig. 10, histograms are color-scaled based on particle type, denoted as $t_{\textrm{ID}}$. Conversely, in Figs. 11 and  12, the color scaling corresponds to gas pressure, $P_{\textrm{gas}}$, and point resolution $\sigma_{r\phi}=\sigma_{z}$ respectively. The CKF results show overall good agreement with the theoretical expectation, with ratios $\sim 1$ for momenta down to 20 MeV/$c$ within statistical uncertainties.

### v2

@@ Abstract
This study introduces a Kalman Filter tailored for homogeneous gas Time Projection Chambers (TPCs), adapted from the algorithm utilized by the ALICE experiment. In order to describe semi-circular paths in the plane perpendicular to the magnetic field, we introduce a novel mirror rotation technique into the Kalman Filter algorithm, enabling effective tracking of trajectories of varying lengths, including those with multiple circular paths within the detector, also known as “loopers”. Demonstrated relative improvements of up to 80% in electron momentum resolution and up to 50% in muon and pion momentum resolution underscore the significance of this enhancement. Significant improvements in the reconstruction efficiency for relatively short low momentum “looper” tracks are also shown. Such advancements hold promise not only for the future of the ALICE TPC but also for neutrino high-pressure gas TPCs, where loopers become significant owing to the randomness of production points and their relatively low energies in neutrino interactions. In particular, an improvement in low energy electron reconstruction, for which the production of “looping” tracks is likely and the impact of the new algorithm is directly demonstrated, could significantly impact the quality of flux determination, which in accelerator neutrino experiments relies on the measurement of $\nu_{e}$ electron scatterings.
@@ 2 The Kalman Filter
    Update: Produce an updated estimate for the state vector and the covariance matrix, $s_{k}$ and $C_{k}$, respectively, using information from the current measurement. The update is performed requiring that the covariance for the new estimate is minimized. This is done through the use of the so called Kalman gain, which is defined as:
@@
    The update operation, also known as filtering, then follows as:
@@ 2.1 The custom Kalman Filter
Multiple scattering is treated through the noise correction matrix, $Q_{k}$. At each step the scattering angle can be treated as emerging from a Gaussian distribution with a root mean square equal to the Molière angle $\theta_{\textrm{M}}$, which is calculated using the formula given by Lynch and Dahl [37]:
@@
Each step in the evolution of the Kalman Filter can potentially fail, in which case the algorithm is stopped. This can happen mainly in two scenarios: $\sin\phi$ can be calculated to be out of range, i.e. $|\sin\phi|>(1-10^{-7})$ or the particle can lose all its remaining energy. Once the Kalman Filter is stopped, the information for each of the reconstructed points is saved. Flags are used to preserve information on which of the reconstruction steps have been successful and which have failed.
@@
Figure 3: Flow chart describing the CKF algorithm. A seeding algorithm is used to obtain an estimate for the status of the system at the start of the trajectory $k=0$. Energy loss and multiple scattering corrections are applied to the estimate. The fit is then moved to the next point $k\rightarrow k+1$ either by applying the the BKF procedure or by using the mirror rotation, in the case that the limits of the $\sin\phi$ range have been surpassed. The algorithm is iterated point by point until the end of the trajectory is reached. Other minor modifications have been made to the CKF algorithm compared to the BKF in order to make the mirroring operation more stable.
@@ 3 Toy Monte Carlo Simulation
To generate particle samples and validate the CKF algorithm, we employed a toy Monte Carlo (MC) tool called fastMCKalman [31]. This tool, stemming from the AliExternalTrackParam framework in the AliRoot code-base [28], was designed to be complemented by RootInteractive [32], an advanced statistical analysis tool. fastMCKalman has been developed with several objectives in mind: conducting rapid Monte Carlo (MC) simulations to evaluate tracking performance metrics, particle identification, and time-of-flight measurements across various detector setups. It was also designed to facilitate detailed studies on signal distortion in the ALICE detector and the derivation of performance metrics for its Run-3 upgrade and future iterations. In this discussion, we highlight the effectiveness of fastMCKalman in rapid simulation and tracking performance assessment within a TPC setting. It is important to note that this article represents the first use of this fast simulation tool and serves as part of its validation process. The use of fast Monte Carlo tools to complement more extensive generators such as Geant4 [40, 41, 42], are commonplace in high energy physics. The ALICE experiment, for example, implements independent Monte Carlo components in its simulation pipeline, to model detector effects that would be difficult to simulate using a more standard MC generator [43]. This is the envisioned future use of fastMCKalman. However, as of now, only basic effects such as multiple Coulomb scattering and energy loss through ionization have been implemented. In order to ensure the agreement of fastMCKalman with more traditional MC generators, the formulas used to simulate these effects are the same implemented by Geant4. Additionally the agreement of the fastMCKalman simulation with theoretical expectations is tested as part of the validation of the CKF algorithm in Sec. 3.2.
@@
The propagation of the state vector is obtained by applying Eq. 15. At each step the energy loss is calculated using the Bethe-Bloch formula as described in Eq. 16 and multiplying by the distance traveled and the material density. The energy loss is then converted into a $q/{p_{\text{T}}}$ multiplicative factor described in Eq. 15 and smeared with a Landau distribution having a width equal to the $c_{k}$ factor described in Eq. 17. The multiple scattering effects are simulated by calculating the scattering angle distribution root mean square $\theta_{\textrm{M}}$ and from that the process noise matrix $Q$. The diagonal elements of the matrix are then used as the widths of smearing Gaussian distributions that are applied to parameters $s_{2}=\sin\phi,s_{3}=\tan\lambda$ and $s_{4}=q/p_{\text{T}}$. To reproduce the measurement noise encapsulated in matrix $R$, a Gaussian smearing is applied to the position parameters $s_{0}=y$ and $s_{1}=z$. The widths of the distributions are equal to the position resolutions $\sigma_{r\phi}$ and $\sigma_{z}$ respectively.
The propagation continues until any of the following happens: the particle reaches the edges of the detector cylinder; the particle looses all its remaining energy; the track has traversed a predefined maximum number of points; one of the propagation steps fails. Once the track is fully generated, the track fit is done as described in Sec. 2.1. No element of track formation or particle identification is included. After saving all information, a new simulation and track fit start.
@@ 3.1 Sample definition
The tracking pad response as well as the gas properties of the detector are sampled in each simulated event: the resolutions $\sigma_{r\phi}=\sigma_{z}$ are uniformly distributed between 0.1 cm and 0.5 cm and the pressure $P_{\textrm{gas}}$ was randomized between 0.1 atm and 10 atm. The gas composition was taken to be the Ne/CO2/N2 (90/10/5) gas mixture used by the ALICE experiment during Run-1 [6]. The radiation length and density of the gas at atmospheric pressure are $X_{0}=1.2763\times 10^{4}\ \text{cm}$ and $\rho=0.0016265\ \text{g/cm}^{3}$. The particles produced are equally divided in electrons, muons, pions, kaons and protons, corresponding to the ALICE’s convention for particle types, $t_{\textrm{ID}}$, 0, 1, 2, 3, and 4, respectively. The angles $\phi$ and $\lambda$ are fully randomized. The initial $p_{\textrm{T}}$ is sampled from a two-component distribution: a high-$p_{\text{T}}$ component uniformly distributed in $[0,20]$ GeV/c, which covers 70% of the total, and a low-$p_{\text{T}}$ component flat in $1/p_{\textrm{T}}$:
@@
where $p_{\textrm{T}_{\textrm{min}}}=0.01$ GeV$/c$, $p_{\textrm{T}_{\textrm{max}}}=20\ \text{GeV}/c$ and $j$ is a random variable uniformly distributed between 0 and 1. Some key properties of the tracks composing the PS sample are plotted in Fig. 5. These include the $p_{\textrm{T}}$ spectrum, the lever arm $L_{\textrm{Arm}}$ and the number of points per track $N$ separated between primaries and secondaries. The lever arm is defined as the distance in the $xy$ plane between the first and last point in the track. All the tracks included in this and future plots have been successfully reconstructed, unless stated otherwise.
@@
Figure 5: Distributions of (a) transverse momentum $p_{\textrm{T}}$, (b) lever arm $L_{\textrm{Arm}}$ and (c) number of points per track $N$ in the PS sample. In all the plots the distributions for primary and secondary particles are shown separately. The low $p_{\textrm{T}}$ portion of the spectrum is produced via Eq. 28. The spikes around $N=250$ and $L_{\textrm{Arm}}=250$ in the primary sample are explained by the simulated geometry having 250 radial pad layers and the particles starting from the center of the detector.
Primaries and secondaries have analogous $p_{\textrm{T}}$, with the only variations arising from the fact that primaries and secondaries have different reconstruction efficiencies. In particular the reconstruction for low momentum secondaries is more likely to fail than for primaries, resulting in a harder spectrum. This arises from the combination of two factors: the low momentum secondaries have a higher risk of having the start of their track at a $\sin\phi$ angle that is close to the edges of its range, making the application of the mirroring algorithm difficult since not many points are remaining; additionally secondaries are more likely to produce short tracks overall, due to the randomness of their starting points. This is a difficulty which could be at least partially solved, by placing the rotation center in front of each secondary track, to mimic the geometrical construction of the primaries. This solution was however not tested for this study and goes beyond the scope of this work. More significant differences appear in the $L_{\textrm{Arm}}$ and $N$ distributions. Since the primaries all start at the center of the detector, most tracks will cross the detector exiting from the cylinder’s barrel, producing a track with as many points as the 250 pad layers. Alternatively the track can exit from the sides of the detector, producing tracks with a slightly smaller $N$ or be stopped inside the detector having $N<50$. For the secondaries the spread is much more homogeneous and the chance of producing tracks with $N>250$ is more significant.
@@
A second sample containing a total of $10^{5}$ particle tracks was produced to recreate conditions analogous to the ones that would be experienced by a HPgTPC in a accelerator neutrino experiment, such as the ND-GAr detector. The goal for this second sample is to explore the potential performance of such a detector, using realistic particle spectra and spatial resolutions. This sample will be referred to as the high-pressure sample or HP sample. The HP sample is produced with randomized starting positions in the same manner as the ones applied to the secondaries in the PS sample. This is done to emulate the randomness of particle track formation in a neutrino experiment. The detector characteristics are fixed, having the same cylinder dimensions and pad distribution of the previous sample. The point resolutions are taken as $\sigma_{r\phi}=\sigma_{z}=0.1\ \text{cm}$, comparable to what is quoted by ALICE [15]. This figure is also used as the benchmark to estimate the point resolution for ND-GAr, given that the MWPC used by the two detector are going to be the same. The gas is a mixture of argon and methane at a 90 to 10 ratio at 10 atm of pressure, which is the nominal gas suggested for the ND-GAr detector in the DUNE Near Detector [13]. This composition corresponds to a $X_{0}=1.193\times 10^{3}\ \text{cm}$ and a density of $\rho=0.01677\ \text{g/cm}^{3}$. Only three particle types are considered in this case: muons, pions and protons. These were chosen because they are the key particles produced in $\nu_{\mu}$ charged-current interactions that are the most relevant in an accelerator neutrino experiment such as DUNE. The initial transverse momenta are randomized to be uniformly distributed between 0.01 GeV$/c$ and 5 GeV$/c$ and the angles are randomized over the whole spectrum. The $p_{\text{T}}$, $L_{\textrm{Arm}}$ and $N$ distributions for the sample are shown in Fig. 6.
@@ 3.2 Tests and results: parameter scan sample
The pulls were tested for the sample, both for the results of the Seed and for the estimates evaluated at the start of the track, after the full propagation of the CKF: the resulting distributions for all the state vector parameters are shown in Figs. 7 and 8, respectively. All the pull distributions were fitted to a standard Gaussian distribution and were found to be centered in 0 and have $\sigma\sim 1$; this implies that the diagonal elements of the covariance matrices well describe the uncertainties.The only significant deviations can be seen for $s_{4}=q/p_{\text{T}}$ for which $\sigma\sim 1.1$ and to a lesser extent $s_{3}=\tan\lambda$. In both cases the underestimations of the matrix elements are likely due to the approximations made during the energy loss correction step. Due to the fact that the $\text{d}E/\text{d}x$ depends on the momentum of the particle and indirectly its energy, the correction procedure is estimated using a step integration. However this method can only in part compensate for the non-linearity in the energy loss and is not completely reliable over long steps. For this reason the more significant deviations are seen for the Seed algorithm.
@@
Figure 8: Pull distributions obtained after the full propagation of the CKF algorithm over the whole PS sample. All distributions were fitted to a Gaussian function. Results for parameters $s_{0}$ to $s_{4}$ (i.e. $y$, $x$, $\sin\phi$, $\tan\lambda$ and $q/p_{\text{T}}$) are shown from left to right and labeled from (a) to (e) accordingly.
@@
where in our case, $\mu$ corresponds to the true value of state vector, $s^{\text{true}}$, $s$ and $C$ are the estimates obtained from the reconstruction and $n=5$. The Mahalanobis distance, $M$, of a set of points belonging to the distribution $D$, follows a $\chi^{2}$ distribution with $n$ degrees of freedom. One can check if $C$ is well defined, by verifying that the corresponding $M$ follow a $\chi^{2}$ distribution with the correct number of degrees of freedom. In Fig. 9, we show the results of a $\chi^{2}$ fit over the $M$ distribution for the whole sample. The plot on the left shows the results obtained from the Seed algorithm, while the one on the right shows the results after the full CKF propagation. In both cases the n.d.f. obtained with the $\chi^{2}$ fit are very close to $n=5$: both Seed and CKF estimates accurately the covariances of the reconstructed track parameters in the state vector. The only deviations are caused by the same $\text{d}E/\text{d}x$ related issue which was pointed out for the Pull test results. It is also important to note that this test is particularly demanding, since all deviations present in the matrix add up to a single figure of merit.
Figure 9: Mahalanobis distance $M$ distribution for the PS Sample fitted by a standard $\chi^{2}$ p.d.f. showing the results for the $n.d.f.$ parameter. The expected result for a 5-dimensional matrix is $n.d.f.=5$. The results for the Seed and for the fully propagated CKF are shown in plots (a) and (b) respectively.
@@
where $l$ is the length of the track in the $xy$-plane. Note that the value of $1/\left(\beta p_{\text{T}}\right)$ is averaged along the trajectory to take into account energy loss. In Figs. 1011 and 12, the upper plots show the CKF covariance estimates, $\sqrt{C_{44}^{\textrm{CKF}}}$, while the bottom plots show their ratios to the theoretical expectations, $\sqrt{C_{44}^{\textrm{theo}}}$. The upper plots are shown to demonstrate how widely the values of $\sqrt{C_{44}^{\textrm{CKF}}}$ vary and to highlight, in comparison, how stable the ratio with the theoretical expectation is. The points analyzed are randomly taken along the reconstructed tracks and down-sampled to $10\%$ of the total to avoid correlations.
@@
Figure 10: (a) CKF $q/p_{\text{T}}$ resolution $\sigma_{\text{CKF}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{CKF}}}$ as a function of the true $q/p_{\text{T}}$. (b) Ratio of the CKF $q/p_{\text{T}}$ resolution, over the theoretical expectations $\sigma_{\text{theo}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{theo}}}$, as a function of the true $q/p_{\text{T}}$. The histograms include all particles in the PS sample and are color-coded according to their particle types $t_{\textrm{ID}}=(0,1,2,3,4)=(\textrm{e},\mu,\pi,K,\textrm{p})$. Only tracks with a minimum of 10 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical. See Fig. 21 in Appendix for a zoom-in version.
Figure 11: (a) CKF $q/p_{\text{T}}$ resolution $\sigma_{\text{CKF}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{CKF}}}$ as a function of the true $q/p_{\text{T}}$. (b) Ratio of the CKF $q/p_{\text{T}}$ resolution, over the theoretical expectations $\sigma_{\text{theo}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{theo}}}$, as a function of the true $q/p_{\text{T}}$. The histograms include all particles in the PS sample and are color-coded according to the gas pressure $P_{\text{gas}}$ used in the simulation. Only tracks with a minimum of 10 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical.
Figure 12: (a) CKF $q/p_{\text{T}}$ resolution $\sigma_{\text{CKF}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{CKF}}}$ as a function of the true $q/p_{\text{T}}$. (b) Ratio of the CKF $q/p_{\text{T}}$ resolution, over the theoretical expectations $\sigma_{\text{theo}}(q/p_{\text{T}})=\sqrt{C_{44}^{\textrm{theo}}}$, as a function of the true $q/p_{\text{T}}$. The histograms include all particles in the PS sample and are color-coded according to the radial resoulution $\sigma_{r\phi}$ used in the simulation. Only tracks with a minimum of 10 points are considered. These plots have been produced using the interactive analytical tool ROOTInteractive [32]. The error bars are statistical.
@@
In Fig. 10,the histograms are color-scaled based on particle type. Conversely, in Figs. 11 and  12, the color scaling corresponds to gas pressure, $P_{\textrm{gas}}$, and point resolution $\sigma_{r\phi}=\sigma_{z}$ respectively. The CKF results show overall good agreement with the theoretical expectation, with ratios $\sim 1$ for momenta down to 20 MeV/$c$ within statistical uncertainties. In Fig. 21 in the Appendix we present an alternative version of Fig. 10 focusing on the central region $q/p_{\text{T}}\in[-5,5]\ c/\text{GeV}$ which is more statistically significant.